# Accuracy of Discretely Sampled Stochastic Policies in Continuous-time Reinforcement Learning

### v1

@@ 7. Proofs of Main Results
Throughout this proof, we denote by $C\geq 0$ a generic constant independent of $t_{i}$, which may take a different value at each occurrence. Note that $X^{\mathscr{G}}_{0}$ is deterministic and hence is $p$-th power integrable for all $p\geq 2$. Suppose that for some $i=0,\ldots,n-2$, (3.3) admits a unique strong solution $X^{\mathscr{G}}$ on $[0,t_{i}]$ satisfying ${\mathbb{E}}[\sup_{t\in[0,t_{i}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. Observe that by (H.1) and the Cauchy-Schwarz inequality, for all $p\geq 2$,
@@
Since $x\mapsto b(t,x,a)$ and $x\mapsto\sigma(t,x,a)$ are Lipschitz continuous uniformly in $(t,a)$, by [40, Theorems 3.3.1 and 3.4.3], (3.3) with the initial state $X^{\mathscr{G}}_{t_{i}}$ at $t_{i}$ admits a unique strong solution $X^{\mathscr{G}}$ on $[t_{i},t_{i+1}]$ satisfying ${\mathbb{E}}[\sup_{t\in[t_{i},t_{i+1}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. This along with an inductive argument shows the desired well-posedness result.
@@
We first prove $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. Observe that by (H.5), for all $p\geq 2$,
@@
Note that $\mathbb{E}^{W}\left[\cdot\right]=\mathbb{E}[\cdot|\widetilde{\mathcal{F}}_{0}]$, and all Itô integrals appear here equal to 0 when taking the conditional expectation $\mathbb{E}[\cdot|\widetilde{\mathcal{F}}_{0}]$. Analogous to the proofs of Lemma 3.1 and [40, Theorem 3.4.3], we obtain that $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$.
Next, we aim to prove that the moments of $X^{\mathscr{G}}_{t}$ is independent of $|\mathscr{G}|$ and obtain (7.10). Use the notation $\delta(u):=t_{i}$ for $u\in[t_{i},t_{i+1})$, then it satisfies that

### v2

@@ Acknowledgments
Throughout this proof, we denote by $C\geq 0$ a generic constant independent of $t_{i}$, which may take a different value at each occurrence. Note that $X^{\mathscr{G}}_{0}$ is deterministic and hence is $p$-th power integrable for all $p\geq 2$. Suppose that for some $i=0,\ldots,n-2$, (3.3) admits a unique strong solution $X^{\mathscr{G}}$ on $0,t_{i}]$ satisfying ${\mathbb{E}}[\sup_{t\in[0,t_{i}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. Observe that by (H.1) and the Cauchy-Schwarz inequality, for all $p\geq 2$,
@@
Since $x\mapsto b(t,x,a)$ and $x\mapsto\sigma(t,x,a)$ are Lipschitz continuous uniformly in $(t,a)$, by [46, Theorems 3.3.1 and 3.4.3], (3.3) with the initial state $X^{\mathscr{G}}_{t_{i}}$ at $t_{i}$ admits a unique strong solution $X^{\mathscr{G}}$ on $t_{i},t_{i+1}]$ satisfying ${\mathbb{E}}[\sup_{t\in[t_{i},t_{i+1}]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. This along with an inductive argument shows the desired well-posedness result.
@@
Assume that $s\in[t_{i-1},t_{i}]$, then $\delta(s)=t_{i-1}$. By Itô’s formula,
$\mathbb{E}[g(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-g(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i_{1}}})]$ (A.5)
$=\mathbb{E}\bigg[\int_{t_{i-1}}^{s}\left(\partial_{t}g(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})+b^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}g(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\right)\mathrm{d}u$
$\quad+\int_{t_{i-1}}^{s}\frac{1}{2}\operatorname*{Tr}\left(\sigma\sigma^{\top}(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\partial_{x}^{2}g(u,X^{\mathscr{G}}_{u},a_{t_{i-1}})\right)\mathrm{d}u\bigg].$
It follows from the linear growth condition of $b$ and $\sigma$ that
$\left|\mathbb{E}[g(s,X^{\mathscr{G}}_{s},a_{t_{i-1}})-g(t_{i-1},X^{\mathscr{G}}_{t_{i-1}},a_{t_{i_{1}}})]\right|$ (A.6)
$\leq\int_{t_{i-1}}^{s}C\|g\|_{C_{p}^{2,0}}\mathbb{E}\big[(1+|X^{\mathscr{G}}_{u}|^{p}+d_{A}(a_{t_{i-1}},a_{0})^{p})\big(1+(1+|X^{\mathscr{G}}_{u}|+d_{A}(a_{t_{i-1}},a))$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{2}(1+|X^{\mathscr{G}}_{u}|+d_{A}(a_{t_{i-1}},a))^{2}\big)\big]\mathrm{d}u$
$\leq C_{p}\|g\|_{C_{p}^{2,0}}|\mathscr{G}|,$
for a constant $C_{p}$ depending only on $b,\sigma$ and $p$, where the last inequality used Lemma 3.1. ∎
Proof of Lemma 6.3.
We first prove $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$. By (H.5), for all $p\geq 2$,
@@
Note that $\mathbb{E}^{W}\left[\cdot\right]=\mathbb{E}[\cdot|\widetilde{\mathcal{F}}_{0}]$, and all Itô integrals appear here equal to 0 when taking the conditional expectation $\mathbb{E}[\cdot|\widetilde{\mathcal{F}}_{0}]$. Analogous to the proofs of Lemma 3.1 and [46, Theorem 3.4.3], we obtain that $\mathbb{E}^{W}[\sup_{t\in[0,T]}|X^{\mathscr{G}}_{t}|^{p}]<\infty$ for all $p\geq 2$.
Next, we prove that the moments of $X^{\mathscr{G}}_{t}$ is independent of $|\mathscr{G}|$ and obtain (6.10). Use the notation $\delta(u):=t_{i}$ for $u\in[t_{i},t_{i+1})$, then it satisfies that